# Constructions are Revealed in Word Distributions

### v1

@@ D.1 Data
Idiomatic usage: Jay comes from a different end of the spectrum to Dave Ambrose, but the two both like to talk nuts and bolts.
@@ D.2 Methods
For each sentence in the dataset we calculate the global affinity for each of the words that are part of the potentially idiomatic expression (PIE). A receiver operating characteristic curve is produced for all words that are part of any PIE using global affinity as classification signal. The plot in the main text includes only examples in which the example sentence is at least ten words long (sufficient context), and where the individual word is at least four characters (avoiding, e.g., determiners and other words). If no filtering is performed the AUC for the ROC is 0.69 (versus 0.71).
Figure 13: ROC curve reflects informativeness of global affinity in figurative vs literal usages of PIEs.
Appendix E Supplement: NPNs (Section 6.1)
Figure 14: Same as Figure 6, but with all sentences shown. The distributions are more skewed to lower probabilities: less acceptable sentences result in lower global affinities.
@@ E.1 Methods
As described in the main text, we filter to generations where the model correctly produces the desired NPN combination in the form noun+prep+noun, resulting in 295 sentences (upon: 97, after: 81, by: 97, to: 20).
To obtain acceptability judgements, we randomly sort all 295 sentences and the last author annotates with a score between 1 and 5, inclusive. When filtered to sentences with acceptabilities of at least 4, we have 162 sentences (upon: 64, after: 59, by: 32, to: 7) and 324 nouns (two for each sentence).
E.2 Data
We provide selected examples from our results:
@@ E.2 Data
    RoBERTa sometimes produces a distribution that favors a date for the second noun when the preposition is after:
    GPT4 generation: The factory evolved, implementing automation after automation, becoming a symbol of modern industry.
    Global affinity scores for automation are 98% and 1.7% respectively. The sentence has an acceptability of 5. RoBERTa prefers WWII, 1950, 1945, modernization as completions in the second noun slot. A number of other sentences with low-scoring nouns, after as preposition, and high (4,5) acceptability scores also exhibit this preference for dates in the second slot.
@@
    A low acceptability (score 1) example:
    The author, lost in his craft, scribbled away pen to pen, endlessly working on his novel.
    Global affinity scores are 19% and 0.03%.
@@
    A case where model assigns a low score to an acceptable (score 5) example:
    The film industry, obsessed with success and failure, measures a movie’s performance rating by rating, scrutinizing every detail for potential improvements.
    Global affinity scores are 65% and 1.5%.
Appendix F Supplement: Comparative Correlative (Section 6.2)
To calculate the percentage of the output distribution nucleus that is a comparative adj/adv, we order the outputs by probability and iterate through them until reaching a total probability mass of $p\geq 0.98$ (a nucleus using 0.98). For each sampled word, we substitute it into the original sentence and use Spacy to check whether it is a comparative adverb or comparative adjective. The final score is the proportion of the sample (the 98% nucleus) that is a comparative adjective or adverb.
Of the 108 ($=54\times 2$) candidate slots, 99 of them are singly tokenized and thus amenable to study using our methods.
Appendix G Supplement: Section 7: The Limits of Distributional Analysis
Figure 15: Corresponding global affinity plot for the local affinity plot in Figure 6. hardware has global affinity (probability) of 0.92, and is mostly affected by hammer, a contextual rather than constructional constraint. Discussion in Section 7.
@@ Appendix G Supplement: Section 7: The Limits of Distributional Analysis
Figure 17: Idioms are not always activated by context. kicked the bucket does not have high global affinity (probability between 0 and 1) until it is clear that the old man has died. “….” indicates omission; “—” indicates empty slots at the end of the sentence.
@@ Appendix H Use of AI Assistant
ChatGPT4o was used to produce initial versions of python matplot generation code in some cases. Any code produced was subsequently adapted/ modified. ChatGPT4o was not used to write any part of this paper. Some potential citations were found using ChatGPT4o as a search tool.

### v2

@@ D.1 Data Sample
Figurative usage: Jay comes from a different end of the spectrum to Dave Ambrose, but the two both like to talk nuts and bolts.
@@ D.2 Methods
Each of the 49,395 sentences in MAGPIE has a PIE that is labeled as either figurative or literal. The words that participate in the PIE are annotated with character spans.
We omit 3,944 sentences for which annotation confidence (figurative or literal) is < 99%. We omit 2,016 where labeled word offsets are wrong (i.e. the indicated word does not match the characters in the span). This gives us 45,450 sentences and 117,385 individual PIE word spans (roughly 2.6 words per PIE) for the analysis. Of these, 10,313 sentences (23,484 spans) are literal and 34,138 (95,917) are figurative. Using RoBERTa, 3,556 words were multitokenized and there were 39 other errors. These were omitted. Our analysis is conducted on the remaining 113,790 singly-tokenized words.
For each labeled character span that is part of a PIE, we simply calculate the global affinity and treat it as a classification signal for figurative versus literal use. Our reported AUC of 0.71 in the main text includes only dataset entries in which the example sentence is at least ten words long (sufficient context), and where the individual word is at least four characters (avoiding, e.g., determiners and other short words). If no filtering is performed the AUC for the ROC is 0.69 (versus 0.71).
Figure 15: ROC curve reflects informativeness of global affinity in figurative vs. literal usages of PIEs.
D.3 Figurative vs. Literal Analysis
Figure 22 compares per-idiom figurative versus literal averages global affinity scores. Only idioms with at least 5 example sentences for both figurative and literal are shown (203 total). Idioms are sorted by figurative score. Brackets give number of examples of each type: [#figurative, #literal]. Idioms where figurative (green) is higher than literal (orange) suggest “success” of the method (e.g., in a rut, nuts and bolts).
Though we do not conduct a full analysis of MAGPIE in this study, we report a few examples from qualitative analysis. In the same way that affinity was able to help identify mislabeled examples in the CEC dataset, affinity draws attention to certain issues or areas of interest in MAGPIE.
For example, consider join the club (0.20 fig vs. 0.49 lit), which we examine to understand why the literal score is higher than the figurative. MAGPIE’s literal examples include usages of the form join a club (e.g., A player joining a new club…). In figurative usages, join the club does not generally admit of lexical or syntactic modification. High affinity for literals reflects contextual activation of join a club or join a/an X club type usages. Moreover, of the few (6) figurative usages in the dataset, many are quoted discourse (“Join the club,” said Connie), which do not produce sufficient activation since other completions like “Join the group” would be valid (hence low affinities). In this dataset, join a/an X club literal usages tend to be longer and better formed. A fairer comparison might enforce a common context length.
Similarly, for play the field (0.22 fig, 0.42 lit; figuratively meaning to hold an interest in a number of people or things), many literal usages refer to a playing field (for sports). The syntactic patterns clearly differ, and affinity has no mechanism to attend to this. For example, playing fields in playing fields and football pitches is entrenched as its own collocation that is likely unrelated to the entrenchment of the figurative play the field.
Sometimes low scoring figurative usages or high scoring literal usages are mislabeled. For example, the top scoring literal example for in black and white (0.92 fig, 0.90 lit overall) is actually mislabeled as literal: …but the complicated plot is hard to follow and the characters are starkly drawn in black and white.
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Figure 16: Same as Figure 7, but with all sentences shown (no acceptability filter). Scores skew lower, reflecting that less acceptable sentences have lower global affinities.
Figure 17: Global affinity for NPNs using the challenge dataset of entirely unattested NPNs with acceptability $\geq 4$. Compare to the non-challenge result shown in Figure 7.
Figure 18: Same as Figure 18 but with all sentences shown (no acceptability filter).
Appendix E Supplement for NPNs (Section 6.1)
@@ E.1 Methods
We verify that generations match the desired form noun+prep+noun.
To obtain acceptability judgements, we randomly sort all sentences and the last author annotates with a score between 1 and 5, inclusive. During acceptability judgement, 2 of the 100 sampled nouns were deemed to be inappropriate, and thus we omitted (2 $\times$ 4 = 8) of the generations.
We produce two datasets: The first is a random sample from all of RoBERTa’s singly-tokenized, singular nouns. The second uses the same sampling procedure but rejects any sampled noun for which the infinigram API has a non-zero count in Pile-train for any of the four NPNs (i.e., the sampled noun with any of the four prepositions).
E.2 Infinigram and the Pile
RoBERTa is trained on five corpora and comprises 160GB of text. Rather than recreating the RoBERTa dataset for our frequency analysis, we search the Pile-train (800GB) using the infinigram API. The Pile is roughly five times larger than RoBERTa’s training data and the five datasets on which RoBERTa was trained are likely included, or partially included, in the Pile:
  • A newer version of BookCorpus is included in Pile
  • English wiki is included in the Pile
  • CC-news is likely included in the Pile-CC, and we verify this by spot-checking CC-news examples using infinigram, finding that most queries are successful
  • A newer version of OpenWebText is included
  • Stories, as part of CommonCrawl is likely included in the Pile-CC, though spot-checks do not find all queried strings
E.3 NPN Dataset: Random sample
When filtered to sentences with acceptabilities of at least 4, we have 244 sentences (upon: 65, after: 73, by: 52, to: 54) and 488 nouns (two for each sentence).
E.4 NPN Dataset: Zero frequency in the Pile
We initially sample 94 nouns and censor 3 of them giving 91 nouns. We generate 364 sentences. When filtered to sentences with acceptability of at least 4, we have 171 sentences (upon: 51, after: 54, by: 34, to: 32).
Figure 18 shows that affinities are lower for this challenge set, but most NPNs using upon still have high affinities, and over half of NPNs using after have high affinity (median affinity $\geq 0.9$. This provides good evidence for generalization of the NPN with these prepositions. NPNs using by and to have lower affinity scores, which accords with the view that they are less productive. Again we observe lower overall affinities when we include sentences judged to be unacceptable (Figure 18).
E.5 Dataset examples
We provide selected examples from our dataset:
@@ E.5 Dataset examples
    Generation in the random sample with high acceptability:
    As a diligent scholar, he poured over his research, analyzing manuscript after manuscript to ensure the accuracy of his findings.
    Acceptability 5; affinity scores: 99.7, 99.7.
@@
    Generation in random sample, with low acceptability:
    Through the philosophical discussions, they delved deeper into the subject, unraveling ambiguity by ambiguity, until clarity was achieved.
    Acceptability 1; affinity scores 99.1, 96.2.
@@
    Generation from the challenge dataset with high acceptability:
    They lived a nomadic life, moving from resettlement to resettlement, always searching for a place to call home.
    Acceptability 5; affinity scores 98.4, 96.0.
  4. 4.
    Generation from the challenge dataset with low acceptability:
    The two rival politicians went ire to ire in a heated debate.
    Acceptability 1; affinity scores 0.0, 0.0.
Appendix F Supplement for Comparative Correlative (Section 6.2)
For each sampled word, we substitute it into the original sentence and use Spacy to check whether it is a comparative adverb or comparative adjective. To calculate the percentage of the output distribution nucleus that is a comparative adj/adv, we order the outputs by probability and iterate through them until reaching a total probability mass of $p\geq 0.98$ (a nucleus using 0.98). The final score is the proportion of the sample (the 98% nucleus) that is a comparative adjective or adverb.
Of the 108 ($54\times 2$) candidate slots, 99 of them are singly tokenized and thus amenable to study using our methods.
Appendix G Supplement for Section 7: The Limits of Distributional Analysis
Figure 19: Corresponding global affinity plot for the local affinity plot in Figure 7. hardware has global affinity (probability) of 0.92, and is mostly affected by hammer, a contextual rather than constructional constraint. See discussion in §7.
@@ Appendix G Supplement for Section 7: The Limits of Distributional Analysis
Figure 21: Idioms are not always activated by context. kicked the bucket does not have high global affinity until it is clear that the old man has died. (The squares for kicked and bucket are dark only in the final sentence.)
@@ Appendix H Use of AI Assistant
ChatGPT4o was used to produce initial versions of python matplot generation code in some cases. Any code produced was subsequently adapted/ modified. ChatGPT4o was not used to write any part of this paper. Some candidate related work was found using ChatGPT4o as a search tool.
Figure 22: Per-idiom global affinity scores for MAGPIE: average affinity for figurative uses (green) and literal uses (orange). Only idioms with at least 5 example sentences for both figurative and literal are shown (203 total). Idioms are sorted by figurative score. Brackets give number of examples of each type: [#figurative, #literal]. Idioms where figurative (green) is higher than literal (orange) suggest “success” of the method.